1. Even though bond markets are larger than equity markets,bond markets tend to be less liquid than equity markets, Evaluate the words in italics. True or False? rue b. False 2. Which of the following statements is most likely FALSE: a. Increasing the number of stocks in a portfolio from 10 to 20 will decrease total risk b. Increasing the number of stocks in a portfolio from 10 to 20 will decrease non-systematic c. Decreasing the number of stocks in a portfolio from 20 to 10 wil increase non-systematic risk d. Decreasing the number of stocks in a portfolio from 20 to 10 wil increase total risk e. Decreasing the number of stocks in a portfolio from 20 to 10 wil increase systematic risk 3. You have $2000 in a savings account today (T 0). Using the Rule of 72, approximate how much longer would it take for your investment to quadruple (i.e. to become worth $8000, or double two times) if you earned 8.0% on your savings vs. 24.0% on your savings at the end of every year? a. 6 years b. 9 years C. ears ears ears positive and negative future cash flows, Evaluate the underlined words in italics. True or False? Hint Think non-conventional cash flows rue b. False 5. A technical analyst believes a stock should be trading at $90, a fundamental analyst uses public information to arrive at a fair price of $80, and the companys management believes the stock is worth $60. Currently, the stock is trading at $70. If you believe the semi-strong form of market efficiency holds, which of the following actions would make the most sense for you to take with the cash in your brokerage account: a. After reviewing the stock price chart, you agree with the technical analyst, and you buy the stock which is b. After reviewing the publically available financial statements, you agree with the sell-side analyst that the c. After reviewing the publically available financial statements, you agree with the sell-side analyst that the d. After calling your sister who works for the company, you learn that clinical trials for the companys phase e. After calling your sister who works for the company, you learn that clinical trials for the company phase II undervalued so you can profit when the stock price increases company is overvalued, and you decide to sell the stock to profit when the stock price declines company is undervalued, and you decide to purchase the stock to profit when the stock price increases IlI drug are not going well; you hold your cash and wait until this information is released to the public drug the company produces are not going well; you short the stock until this news is released to the public
Answer #1

1. False.

2. Option e is Correct Option. Decreasing number if stocks does not increase systematic risk

3. Quadrupling at 8% = 2* 72/8= 18
Quadrupling at 24% = 2* 72/24= 6

So difference = 18-6 = 12 years

4.False. IRR can give multiple values but NPV always gives a single value.